OpenStudy (anonymous):

A buffer has a pH of 4.87. If the buffer is made from a weak acid (Ka = 3.30 * 10^-5), and its conjugate base, the [conjugate base]/[weak acid] ratio is _______.

OpenStudy (anonymous):

\[4.87=-\log(Ka)+\log (\frac{ base }{ acid })\] and just replace the (bae/acid) as x and it should work out fine

OpenStudy (anonymous):

I got 2.45 just for reference

OpenStudy (anonymous):

Thank you!